OpenStudy (azualvar):

The product of a non-zero rational number and an irrational number can always be written as a fraction. a terminating decimal. a non-repeating, non-terminating decimal. a repeating decimal

OpenStudy (sshayer):

what do you think?

OpenStudy (sshayer):

example \[\frac{ 3 }{ 4 }\times \sqrt{7}=\frac{ 3\sqrt{7} }{ 4 }=?\] \[\sqrt{7}=2.645751311...\]